If the operating potential in an X-ray tube is increased by 1%, by what percentage does the cutoff wavelength is 30 pm, find the electric field between the cathode and the target.

Grade:10

1 Answers

Jitender Pal
9 years ago
we know λ = hc/V Now λ = hc/1.01 = λ/1.01 λ – λ’ = 0.01/1.01 λ. % change of wave length = 0.01* λ/1.01 * λ * 100 = 1/1.01 = 0.9900 = 1%.